# EffoVPR: Effective Foundation Model Utilization for Visual Place Recognition

### v1

@@ Abstract
The task of Visual Place Recognition (VPR) is to predict the location of a query image from a database of geo-tagged images. Recent studies in VPR have highlighted the significant advantage of employing pre-trained foundation models like DINOv2 for the VPR task. However, these models are often deemed inadequate for VPR without further fine-tuning on task-specific data. In this paper, we propose a simple yet powerful approach to better exploit the potential of a foundation model for VPR. We first demonstrate that features extracted from self-attention layers can serve as a powerful re-ranker for VPR. Utilizing these features in a zero-shot manner, our method surpasses previous zero-shot methods and achieves competitive results compared to supervised methods across multiple datasets. Subsequently, we demonstrate that a single-stage method leveraging internal ViT layers for pooling can generate global features that achieve state-of-the-art results, even when reduced to a dimensionality as low as 128D. Nevertheless, incorporating our local foundation features for re-ranking, expands this gap. Our approach further demonstrates remarkable robustness and generalization, achieving state-of-the-art results, with a significant gap, in challenging scenarios, involving occlusion, day-night variations, and seasonal changes.
@@ 1 Introduction
The task of Visual Place Recognition (VPR), also known as Geo-Localization, aims to predict the place where a photo was taken relying solely on the visual information in the image. This is typically done by an image retrieval approach [4, 21, 34, 11, 33, 27] where a database of geo-tagged images is used, often referred as gallery. Real world data, including tagged VPR datasets, rely on two major sources for images: 1) car street-view and 2) people personal cameras (commonly mobile phones) [4, 43, 47, 8, 6]. As a result, images contain natural objects that are irrelevant and sometimes misleading for VPR task. Figure 1 - top row, shows an example, where people, vehicles, daylight, or camera angles might differ between images.
@@
Modern models commonly use a deep neural network to extract a so-called global feature (a.k.a descriptor) for the query and gallery images [29, 24, 39, 30, 31]. This single-stage approach is then followed by a nearest neighbor search in the feature space, to retrieve the matching candidates from the gallery. In practice, global features of the entire gallery need to be uploaded to RAM to enable fast retrieval. In large-scale and real-world scenarios, reducing the memory footprint for each image is crucial to ensure real-time applicability. Therefore, several works often promote compact features to achieve both accuracy and applicability [49, 34, 8]. A popular strategy to improve the accuracy is to conduct a two-stage search, with subsequent similarity search on the top-k ranked results, that re-ranks the first-stage retrieved results. This process is commonly performed by matching local key-points and their corresponding descriptors.
@@
Following best practices in computer vision, VPR methods are often initialized with ImageNet pre-trained weights (e.g. [49, 9]), followed by finetuning on VPR datasets, e.g. MSLS [47], or trained from scratch [46, 21, 4]. Recent advancements in VPR exploit the capabilities of foundation models [27, 22, 34, 33] such as DINOv2 [37], a transformer [44] based model trained by self-supervised learning on a vast amount of data. Recent approaches [34, 33, 22] argue that using vanilla DINOv2 (without fine-tuning) is ineffective. They criticize this approach for failing by capturing dynamic and irrelevant elements (pedestrians or vehicles), thereby diverting attention from crucial VPR features (buildings or scene layout). Some methods propose to address this issue by integrating learned adaptors into the foundation model architecture [34, 33] or changing the standard fine-tuning scheme with a unique pooling mechanism [22], in order to facilitate the foundation model adaptation to the VPR task. In contrast, Anyloc [27] suggested to use DINOv2 as a general-purpose feature representation at zero-shot for various localization tasks. However, despite achieving its best performance with the extremely high dimensional features (of size 49K), faces scalability challenges, AnyLoc only demonstrates a modest performance, and struggles in cases where queries exhibit large appearance changes. In this paper, we utilize the internal ViT features to firstly surpasses previous zero-shot approaches and achieve competitive results compared to several VPR trained methods.
@@
Typically, VPR methods [22, 34, 33, 11, 49, 8] adhere to a conventional approach of initially extracting local features from a pretrained model, then using pooling methods such as GeM [38] or NetVLAD [4], to obtain a global feature, for each image. Methods using VLAD[23], such as AnyLoc[27], or NetVLAD[4] for aggregating local features necessitate learning a dictionary for each specific gallery. These models tend to "overfit" to the particular gallery distribution, which hampers their generalization. Additionally, they often require fine-grained clustering, leading to high-dimensional feature vectors. In addition, a recent study[40], highlighted the GeM hyper-parameter sensitivity, spatial information loss, and convergence complications. Alternatively, we suggest to utilize the internal ViT self-attention mechanism, and training with the class $[CLS]$ token, for classification loss. This approach allows implicit aggregation from local features to a global representation and eliminates the need for previously used external special components and learned features, that often yield excessively large feature representation [22, 33, 34] (see Figure 1).
@@
We hereby present an Effective foundation based VPR method called EffoVPR which achieves a superior single-stage performance without the need for additional components. Unlike more intricate DINOv2-based methods with built-in adapter components [34, 33] or special aggregation techniques [22], EffoVPR delivers State-of-The-Art (SoTA) performance across multiple datasets while maintaining compactness, as shown in Figure 1. Our two-stage approach, builds upon local feature matching originated from the internal ViT attention maps, thus further expanding our performance gap. Additionally, EffoVPR demonstrates strong generalization and robustness across various cities and landscapes, effectively handling challenges such as occlusions, time differences and seasonal changes as demonstrated in Figure 1, which illustrates significant occlusion.
@@
    We present a method that effectively leverages a foundation model for the VPR task, exploiting its intermediate attention layers. Leveraging this finding, we propose a zero-shot method that outperforms previous zero-shot approaches while showing comparable results, even with trained VPR methods, on few datasets.
@@
    Our work introduces a novel foundation model based VPR approach that eliminates the need for external aggregation methods or specialized pooling layers (such as NetVLAD or GeM).
@@
    We suggest an effective yet simple re-ranking process based on ViT internal attention layers, that significantly boosts performance.
@@
    Our method achieves state-of-the-art performance across various VPR benchmarks and generalizes to challenging scenarios, as occlusions, day-night variations or seasonal changes.
@@ 2 Related work
Traditional approaches utilized a single-stage approach for Visual Place Recognition (VPR), using SIFT [32] SURF [7], or RootSIFT [25], focused on matching queries to gallery images through the use of image local feature matching. Two-stage methods[21, 46, 49, 34] entail an initial ranking based on global representation similarity, succeeded by re-ranking the top-K retrieved candidates in the second stage, utilizing local features.
First deep learning approaches for the VPR task used CNN [4, 38, 26] that was later replaced with Vision Transformers (ViT) [46, 49, 27, 34, 33, 22]. TransVPR [46] and R2Former [49] suggested the application of transformers for VPR and adopted a two-stage approach that included re-ranking. However, training from scratch or initialized on ImageNet, and lack of effective view-variability in training (in contrast to [11, 33]) has restricted their performance.
A handful of visual foundation models (VFMs) have recently emerged as the backbones for numerous tasks. VFMs like CLIP [39], DINOv2 [37], SAM [28] use ViT that are trained with distinct objectives, exhibiting unique characteristics for various downstream applications often even at zero-shot (without requiring any fine-tuning). Among zero-shot applications, AnyLoc [27] proposed a single-stage solution for VPR, that leveraged a DINOv2 pre-trained model alongside dense local (patch-level) features. By employing VLAD for feature pooling across multiple layers, AnyLoc approach produces an extremely large global representation of $\sim$49K dimensions, subsequently reduced to 512D through PCA whitening, but in expense of lower performance. However, AnyLoc’s VLAD aggregation tends to fail in generalizing to queries with large time-gaps, day vs. night or of different season. In this paper, we suggest a two-stage zero-shot method based on first global ranking, then re-ranking, by matching local descriptors, extracted from the ViT self-attention layers.
On the contrary, SALAD [22] proposed fine-tuning the pre-trained DINOv2 model to improve its performance for VPR task. Their single-stage approach involves pooling local features from the output layer of DINOv2 replacing NetVLAD’s soft-assignment to clusters by an optimal transport methodology. Notably, their most significant results are achieved with large features exceeding 8K dimension, which adversely impacts memory consumption.
Using DINOv2 as their backbone, SelaVPR [34] and CricaVPR [33] took a different path and avoided fine-tuning of the model, claiming for a side-effect of a catastrophic forgetting and performance deterioration. They propose the incorporation of a trainable adapters integrated into DINOv2 architecture. SelaVPR suggested a two-stage methodology that trains global and local features on two separate branches, built above DINOv2’s output tokens, while discarding the standard [CLS] token. Their re-ranking strategy relies on mutual-Nearest Neighbor (mutual-NN) computed over a newly learned patch level (local) features. CricaVPR instead, suggests learning a features pyramid above DINOv2’s output tokens, to learn a special viewpoint robust encoder. Both, SelaVPR and CricaVPR incorporate adapters in the DINOv2 architecture, using GeM pooling for feature aggregation and are trained with a contrastive loss.
@@
Training a VPR model often sourced from Street-View images, introduces a significant challenge in selecting appropriate positive and hard negative images[4, 47, 49, 8]. EigenPlaces [11] introduced a novel training paradigm that organizes the training data into classes, with each class containing multiple viewpoints of the same scene. By employing a classification loss rather than the conventional contrastive loss (used in [34, 22, 33]), the resulting model demonstrated high resilience to varying viewpoints in testing. Hence, we adopt a similar strategy for our training process.
@@
Current VPR methods predominantly, derive a global representation by aggregating local features obtained from the output layer of a trained backbone. This aggregation is often conducted using external learnable pooling techniques, such as VLAD (Vector of Locally Aggregated Descriptors) [23], SPoC (Sum-Pooled Convolutional features) [5], RMAC (Regional Maximum Activation of Convolutions) [42], the known NetVLAD [4] or the nowadays widely-used GeM (Generalized Mean) pooling layer[38, 40], as utilized in various studies [12, 18, 19, 47, 21, 22, 34, 33]. However, NetVLAD is encumbered by high computational costs and dimensional complexities (with $\sim$ 32K dimensional feature)[8], while GeM encounter convergence issues and require hyperparameter tuning [40]. In contrast, we suggest to forgo the explicit and external aggregation approach, replacing it with training of the single [CLS] token in ViT, with classification loss.
In summary, this paper revisits the potential of foundation DINOv2 for VPR task, considering two-stage zero-shot and single as well as two-stage finetuned models. Unlike previous approaches, we propose leveraging the existing internal aggregation layers in DINOv2 without any additional components or adaptors. Notably, our approach achieves SoTA results with compact features, a crucial consideration given realistic memory constraints.
@@ 3 Method
For the VPR task, we explore ViT [17] feature maps as local patch descriptors. In a ViT architecture, an image is split into $p$ non-overlapping patches which are processed into tokens by linearly projecting each patch to a $d$-dimensional space, and adding learned positional embeddings. An additional [CLS] token is inserted to capture global image properties. The set of tokens are then passed through $L$ transformer encoder layers, each consists of normalization layers, Multihead Self-Attention modules, and MLP blocks. In ViT, each patch is directly associated with a set of features, a Key, Query, Value that can be used as patch descriptors. We utilize the self-attention matrices at layer $l$, $\{K_{l},Q_{l},V_{l}\}\in\mathbb{R}^{p\times d}$, with $p$ indicating the number of patches resulting $p+1$ tokens (number of patches in the image plus one added global token) and $d$ standing for the embedding dimension. The self-attention function at layer $l$ is then given by:
$\text{Attention}(Q_{l},K_{l},V_{l})=\text{Softmax}(\frac{K_{l}^{T}Q_{l}}{\sqrt{d}})V_{l}$ (1)
We indicate the key feature of the $[CLS]$ token at layer $l$ by $k_{cls}\in K_{l}$. Each image patch therefore is directly associated with the set of features $\{q_{i}^{l},k_{i}^{l},v_{i},^{l}\}_{i\in{[p+1]}}$ including its query, key and value at each layer $l\in[1,L]$, respectively.
In the global search stage, the aim is to perform an efficient search across a vast corpus of images in a gallery. To this end, a shared embedding space for query and gallery images is typically utilized to identify the most similar images and rank them accordingly. In the re-ranking stage, the top-K nearest neighbors (candidates) are further refined by evaluating the mutual similarity between their local features and those of the query image. Throughout both stages, we use a ViT [17] encoder as our backbone model. Our EffoVPR retrieval method is presented in Figure 2.
Figure 2: An overview of EffoVPR. Left: During inference, we identify the nearest neighbors of the query by using the [CLS] token as the global representation for each image ($v_{g}$). For the second re-ranking stage, we extract (dashed-line) intermediate value features $V$ and filter them (with a predefined threshold $T_{1}$) using a score derived from the partial attention maps $S$. Right: Lastly, we re-rank the top-K candidates from the first stage based on the count of strongly connected mutual nearest neighbors (MNN) with a score exceeding a predefined threshold (denoted as $T_{2}$).

### v2

@@ Abstract
The task of Visual Place Recognition (VPR) is to predict the location of a query image from a database of geo-tagged images. Recent studies in VPR have highlighted the significant advantage of employing pre-trained foundation models like DINOv2 for the VPR task. However, these models are often deemed inadequate for VPR without further fine-tuning on VPR-specific data. In this paper, we present an effective approach to harness the potential of a foundation model for VPR. We show that features extracted from self-attention layers can act as a powerful re-ranker for VPR, even in a zero-shot setting. Our method not only outperforms previous zero-shot approaches but also introduces results competitive with several supervised methods. We then show that a single-stage approach utilizing internal ViT layers for pooling can produce global features that achieve state-of-the-art performance, with impressive feature compactness down to 128D. Moreover, integrating our local foundation features for re-ranking further widens this performance gap. Our method also demonstrates exceptional robustness and generalization, setting new state-of-the-art performance, while handling challenging conditions such as occlusion, day-night transitions, and seasonal variations.
@@ 1 Introduction
The task of Visual Place Recognition (VPR), also known as Geo-Localization, aims to predict the place where a photo was taken relying solely on the visual information in the image. This is typically done by an image retrieval approach (Arandjelovic et al., 2016; Hausler et al., 2021; Lu et al., 2024b; Berton et al., 2023; Lu et al., 2024a; Keetha et al., 2023) where a database of geo-tagged images is used, often referred as gallery. Real world data, including tagged VPR datasets, rely on two major sources for images: 1) car street-view and 2) people personal cameras (commonly mobile phones) (Arandjelovic et al., 2016; Torii et al., 2015; Warburg et al., 2020; Berton et al., 2022a; Barbarani et al., 2023). As a result, images contain natural objects that are irrelevant and sometimes misleading for VPR task. Figure 1 (top-row) shows an example, where people, vehicles, daylight, or camera angles might differ between images.
@@
Compact Features: Modern models commonly use a deep neural network to extract a so-called global feature (a.k.a descriptor) for the query and gallery images (Levy et al., 2023; Jia et al., 2021; Radford et al., 2021; Li et al., 2022b; Liu et al., 2021). This so called single-stage approach is then followed by a nearest neighbor search in the feature space, to retrieve the matching candidates from the gallery. In practice, global features of the entire gallery need to be maintained in the RAM memory to enable fast retrieval. In large-scale and real-world scenarios, reducing the memory footprint for each image is crucial to ensure real-time applicability. Therefore, several works often promote compact features to achieve both accuracy and applicability (Zhu et al., 2023; Lu et al., 2024b; Berton et al., 2022a).
@@
Foundation Models: Following best practices in computer vision, VPR methods are often initialized with ImageNet pre-trained weights (e.g. (Zhu et al., 2023; Berton et al., 2022b)), followed by finetuning on VPR datasets, e.g. MSLS (Warburg et al., 2020), or trained from scratch (Wang et al., 2022; Hausler et al., 2021; Arandjelovic et al., 2016). Recent advancements in VPR exploit the capabilities of foundation models (Keetha et al., 2023; Izquierdo & Civera, 2024; Lu et al., 2024b; a) such as DINOv2 (Oquab et al., 2023), a transformer (Vaswani et al., 2017) based model trained by self-supervised learning. Recent approaches (Lu et al., 2024b; a; Izquierdo & Civera, 2024) argue that using vanilla DINOv2 (without fine-tuning) is ineffective. They criticize this approach for failing by capturing dynamic and irrelevant elements (pedestrians or vehicles). Some methods propose to address this issue by integrating learned adaptors into the foundation model architecture (Lu et al., 2024b; a) or changing the standard fine-tuning scheme with a unique pooling mechanism (Izquierdo & Civera, 2024), in order to facilitate the foundation model adaptation to the VPR task. In contrast, Anyloc (Keetha et al., 2023) suggested to use DINOv2 as a general-purpose feature representation at zero-shot for various localization tasks.
@@
Pooling Methods: Typically, VPR methods (Izquierdo & Civera, 2024; Lu et al., 2024b; a; Berton et al., 2023; Zhu et al., 2023; Berton et al., 2022a) adhere to a conventional approach of initially extracting local features from a pretrained model, then using pooling methods such as GeM (Radenović et al., 2018) or NetVLAD (Arandjelovic et al., 2016), to obtain a global feature. Methods using VLAD (Jégou et al., 2010), such as AnyLoc (Keetha et al., 2023), or NetVLAD (Arandjelovic et al., 2016) for aggregating local features necessitate learning a dictionary for each specific gallery. These models tend to “overfit” to the particular gallery distribution, which hampers their generalization. Additionally, they often require fine-grained clustering, leading to high-dimensional feature vectors. Furthermore Shao et al. (2023), highlighted GeM shortcomings, as hyper-parameter sensitivity, spatial information loss, and convergence complications.
In this paper, we challenge the necessity of the external pooling layers, and suggest to utilize the internal ViT self-attention mechanism for an implicit pooling (Li et al., 2022a). Apparently, a straight-forward approach that trains with the class $[CLS]$ token solely (instead of finetuning by all patch tokens) implicitly aggregates local features into a global representation. This approach removes the reliance on external specialized components or learned features, which often result in excessively large feature representations (Izquierdo & Civera, 2024; Lu et al., 2024a; b; Ali-bey et al., 2024). Moreover, this approach facilitates simple dimensionality reduction, allowing to achieve state-of-the-art results even with compact feature representations (see Figure 1 - bottom row).
@@
Re-ranking: A popular strategy to improve the accuracy is to conduct a two-stage search, with subsequent similarity search on the top-k ranked results. This step is commonly performed by matching local key-points and their corresponding descriptors. We build on prior work highlighting the robust properties of SSL-trained models, e.g. their semantic and instance-level capabilities Oquab et al. (2023); Caron et al. (2021); Park et al. (2023); Keetha et al. (2023), with Amir et al. (2022), exploring the roles of self-attention facets across various tasks. In this paper, we propose a matching method for VPR that uses specific internal ViT features, effective even in a zero-shot setting, outperforming current zero-shot methods and competitive with several fine-tuned VPR approaches. Fine-tuning the DINOv2 backbone for the VPR task further enhances the quality of the features and the overall performance of our matching method, achieving state-of-the-art results.
We hereby present an Effective foundation based VPR method called EffoVPR that achieves superior single-stage performance without requiring additional components. Unlike more intricate DINOv2-based methods (Lu et al., 2024a) that rely on adapter components (Lu et al., 2024b) or special aggregation techniques (Izquierdo & Civera, 2024), EffoVPR delivers state-of-the-art results across multiple datasets while maintaining a compact feature representation. Additionally, EffoVPR demonstrates strong generalization and robustness across diverse cities and landscapes, effectively handling challenges such as occlusions, time differences and seasonal changes as demonstrated in Figure 1, which illustrates significant occlusion.
@@
    We introduce a novel image matching method for VPR that relies on keypoint detection and descriptor extraction from the intermediate layers of a foundation model. This method outperforms previous zero-shot approaches while showing comparable results with trained VPR methods on a few datasets.
@@
    We challenge the necessity of aggregation methods commonly used in various state-of-the-art methods, and propose a simple yet effective alternative that significantly improves performance and feature compactness.
@@
    We propose an effective approach that involves fine-tuning, followed by re-ranking the top candidates using our matching method. This approach significantly boosts performance, even with a very small number of candidates.
@@
    We achieve state-of-the-art results on numerous benchmarks, demonstrating outstanding robustness across varying scenes e.g. day vs. night, different seasons and occlusions, while achieving feature compactness that is often orders of magnitude better than prior methods.
@@ 2 Related work
Traditional approaches utilized a single-stage approach for Visual Place Recognition (VPR), using SIFT (Lowe, 2004) SURF (Bay et al., 2006), or RootSIFT (Jin et al., 2021), focused on matching queries to gallery images through the use of image local feature matching. Two-stage methods (Hausler et al., 2021; Wang et al., 2022; Zhu et al., 2023; Lu et al., 2024b) entail an initial ranking, based on global representation similarity, succeeded by re-ranking the top-K retrieved candidates in the second stage, utilizing local features.
@@
First deep learning approaches for the VPR task used CNN (Arandjelovic et al., 2016; Radenović et al., 2018; Jin Kim et al., 2017) that was later replaced with Vision Transformers (ViT) (Wang et al., 2022; Zhu et al., 2023; Keetha et al., 2023; Lu et al., 2024b; a; Izquierdo & Civera, 2024). TransVPR (Wang et al., 2022) and R2Former (Zhu et al., 2023) suggested the application of transformers for VPR and adopted a two-stage approach that included re-ranking. However, training from scratch or initialized on ImageNet, and lack of effective view-variability in training has restricted their performance as they introduce a significant challenge in selecting appropriate positive and hard negative images (Arandjelovic et al., 2016; Warburg et al., 2020; Zhu et al., 2023; Berton et al., 2022a). EigenPlaces (Berton et al., 2023) introduced a novel training paradigm that organizes the training data into classes, with each class containing multiple viewpoints of the same scene. By employing a classification loss rather than the conventional contrastive loss (used in (Lu et al., 2024b; Izquierdo & Civera, 2024; Lu et al., 2024a)), the resulting model demonstrated high resilience to varying viewpoints in testing. Hence, we adopt a similar strategy for our training process.
A handful of visual foundation models (VFMs) have recently emerged as the backbones for numerous tasks. VFMs like CLIP (Radford et al., 2021), DINOv2 (Oquab et al., 2023), SAM (Kirillov et al., 2023) use ViT that are trained with distinct objectives, exhibiting unique characteristics for various downstream applications. In VPR, SALAD (Izquierdo & Civera, 2024) proposed fine-tuning the pre-trained DINOv2 model to improve its performance for VPR task. Their single-stage approach involves pooling local features from the output layer of DINOv2 replacing NetVLAD’s soft-assignment to clusters by an optimal transport methodology. Notably, their most significant results are achieved with large features exceeding 8K dimension, which adversely impacts memory consumption. Using DINOv2 as their backbone, SelaVPR (Lu et al., 2024b) and CricaVPR (Lu et al., 2024a) took a different path and avoided fine-tuning of the model, proposing the incorporation of a trainable adapters integrated into DINOv2 architecture. SelaVPR built above DINOv2’s output tokens, while discarding the standard [CLS] token. Their re-ranking strategy does not require spatial verification and is based on a newly learned patch level (local) features. CricaVPR instead, suggests learning a features pyramid above DINOv2’s output tokens, to learn a special viewpoint robust encoder. Both, SelaVPR and CricaVPR incorporate adapters in the DINOv2 architecture, using GeM pooling for feature aggregation, trained with the common contrastive loss.
@@
Current VPR methods predominantly, derive a global representation by aggregating local features obtained from the output layer of a trained backbone. This aggregation is often conducted using external modules for learnable pooling techniques, such as VLAD (Vector of Locally Aggregated Descriptors) (Jégou et al., 2010), SPoC (Sum-Pooled Convolutional features) (Babenko & Lempitsky, 2015), RMAC (Regional Maximum Activation of Convolutions) (Tolias et al., 2015), the known NetVLAD (Arandjelovic et al., 2016) or the nowadays widely-used GeM (Generalized Mean) pooling layer (Radenović et al., 2018; Shao et al., 2023), as utilized in various studies (Berton et al., 2021; Garg & Milford, 2021; Ge et al., 2020; Warburg et al., 2020; Hausler et al., 2021; Izquierdo & Civera, 2024; Lu et al., 2024b; a). However, NetVLAD is encumbered by high computational costs and dimensional complexities (with $\sim$ 32K dimensional feature) (Berton et al., 2022a), while GeM encounter convergence issues and require hyperparameter tuning (Shao et al., 2023). Recently, more sophisticated approaches were suggested. SALAD (Izquierdo & Civera, 2024) applied optimal transport for local features aggregation, while BoQ (Ali-bey et al., 2024) aggregated features by learning a Bag-of-Queries along cross-attention. In this paper, we propose a method that, unlike previous approaches, utilizes the existing internal aggregation layer in DINOv2 without the need for additional pooling components or adaptors, achieving SoTA results with highly compact features.
DINOv2 deep features have been utilized for various tasks, e.g. part co-segmentation (Amir et al., 2022), establishing semantic correspondences between image pairs (Zhang et al., 2023; Shtedritski et al., 2023; Zhang et al., 2024) or visual appearance transfer (Tumanyan et al., 2022), using internal features. Motivated by these studies we present a novel local matching method for VPR that detects keypoints and extracts their corresponding descriptors using a unique combination of ViT features.
Foundation models has shown strong applications even at zero-shot setting (without requiring any fine-tuning). In VPR, AnyLoc (Keetha et al., 2023) proposed a single-stage solution that leveraged a DINOv2 pre-trained model alongside dense local (patch-level) features. By employing VLAD for feature pooling across multiple layers, AnyLoc approach produces an extremely large global representation of $\sim$49K dimensions, subsequently reduced to 512D through PCA whitening, but in expense of lower performance. AnyLoc’s VLAD aggregation tends to fail in generalizing to queries with large time-gaps, day vs. night or of different season. In this paper, we propose a two-stage zero-shot approach for VPR that first performs global ranking, followed by re-ranking. Moreover, our approach is adaptable to fine-tuned models. Remarkably, after fine-tuning, it attains state-of-the-art results while preserving compact features, which is essential given practical memory constraints.
@@ 3 Method
We start by outlining our model training strategy, followed by our local feature extraction method. Our approach stems from the insight that DINOv2 foundation model encapsulates robust features suitable for VPR tasks (Keetha et al., 2023). The core idea is to employ specific facets within the self-attention layer as reliable keypoint detectors and others as robust descriptors, for re-ranking.
We explore ViT (Dosovitskiy et al., 2021) feature maps as local patch descriptors. In a ViT architecture, an image is split into $p$ non-overlapping patches which are processed into tokens by linearly projecting each patch to a $d$-dimensional space, and adding learned positional embeddings. An additional [CLS] token is inserted to capture global image properties. The set of tokens are then passed through $L$ transformer encoder layers, each consists of normalization layers, Multihead Self-Attention modules, and MLP blocks. In ViT, each patch is directly associated with a set of features, a Key, Query, Value that can be used as patch descriptors. We utilize the self-attention matrices at layer $l$, as $\{K,Q,V\}\in\mathbb{R}^{(p+1)\times d}$, with $p$ indicating the number of patches resulting $p+1$ tokens (number of patches in the image plus one added global token) and $d$ standing for the embedding dimension. The self-attention function at layer $l$ is then given by:
$\text{Attention}(Q,K,V)=\text{Softmax}(\frac{K^{T}Q}{\sqrt{d}})V$ (1)
Note that for sake of brevity we omit the index $l$. We indicate the key feature of the $[CLS]$ token at layer $l$ by $k_{cls}\in K$. Each image patch therefore is directly associated with the set of features $\{q_{i},k_{i},v_{i}\}_{i\in{[p+1]}}$ including its query, key and value at certain layer $l\in[1,L]$, respectively.
In the global search stage, the aim is to perform an efficient search across a vast corpus of images in a gallery. To this end, a shared embedding space for query and gallery images is typically utilized to identify the most similar images and rank them accordingly. In the re-ranking stage, the top-K nearest neighbors (candidates) are further refined by evaluating the mutual similarity between their local features and those of the query image. Throughout both stages, we use a ViT (Dosovitskiy et al., 2021) encoder as our backbone model. Our EffoVPR retrieval method is presented in Figure 2.
Figure 2: An overview of EffoVPR. Left: During inference, we identify the nearest neighbors of the query by using the [CLS] token as the global representation for each image ($v_{g}$). For the second re-ranking stage, we extract (dashed-line) intermediate features, and utilize $S$, the partial attention map, for keypoint selection (with a predefined threshold $T_{1}$), while employing the Value facet $V$ as the corresponding keypoints descriptors. Right: Lastly, we re-rank the top-K candidates from the first stage based on the count of strongly connected mutual nearest neighbors (MNN) with a score exceeding a predefined threshold (denoted as $T_{2}$).